A child standing on the edge of a sheer cliff wall throws a rock down towards the ground below. The rock is thrown 27° below the horizontal at a speed of 14m/s & lands 13 m from the base of the cliff wall. Ignore air drag.

A.) Determine how long the rock was in the air.


B.) Determine how high the cliff wall is.


C.) Determine how fast the rock was going upon impact.


D.) Determine the angle of impact.
°
Answer #1

a)

Considering horizontal motion

x = u cos x * t

13 = 14* cos 27 * t

t = 1.042 s

=========

b)

Using 2nd equation of motion

h = u sin x * t + 0.5* gt^2

h = 14* sin 27* 1.042 + 4.9 (1.042)^2

h = 11.943 m

=======

C)

Vertical velocity

Vy = u sin x + gt = 14 sin 27 + 9.8* 1.042 = 16.57 m/s

Horizontal velocity

Vx = u cos x = 14 cos 27 = 12.474 m/s

Velocity just before impact

v^2 = vx^2 + vy^2

v = 20.74 m/s

========

D)

Direction,

Phi = arctan (vy/Vx)

Phi = 53 deg below horizontal

========